What would be the dimensions if the area of a table is 1650in and the length is 17 inches more than the width?
Answer by unlockmath(1688)

Hello,
We can use x to show the width and x+17 to be the length. Now we can set up the following equation: (Note: it should be 1650 Square inches)
x(x+17)=1650 Rewritten to look:
x^2+17x-1650=0 Factor this to get:
(x-33)(x+50)=0 Solve x:
x=33
x=-50
The obvious answer is 33 so the width is 33 inches and length is 50 inches.
Make sense?
Check out a book I wrote on the basics of math:
www.math-unlock.com